OpenStudy (anonymous):

simplify the following expression. -4x( -2x^2 +3y -1)

OpenStudy (anonymous):

a(b+c) = ab+ac

OpenStudy (anonymous):

okay den

OpenStudy (anonymous):

use that property to solve your expression

OpenStudy (anonymous):

so -4x(-2x^2+3y+1) = -4x(-2x^2) + -4x(+3y) + -4x(+1)